A cash register contains only five dollar and ten dollar bills. It contains twice as many five’s as ten’s and the total amount of money in the cash register is 600 dollars. How many ten’s are in the cash register?
Answer by richwmiller(17219)

5f+10t=600
f=2t
substitute
5*2t+10*t=600
20t=600
t=30
f=60
check
5*60+10*30=600
300+300=600
ok